OpenStudy (joftheworld):

6x + 3y = 12 6x - 5y = 60

OpenStudy (anonymous):

y=9

OpenStudy (anonymous):

x=-2.5

OpenStudy (joftheworld):

-7x - 6y = 59 -8x -6y = 58